OpenStudy (jiteshmeghwal9):

some knowledge about \(\Huge{\color{red}{\pi}}\) \(\Huge{\color{blue}{\ddot \smile}}\)

OpenStudy (kropot72):

Pi is the ratio of the circumference of a circle to its diameter. Pi is an irrational number meaning that it cannot be expressed as a fraction (an integer over an integer)

OpenStudy (anonymous):

\(\huge \color{red}{\pi} = \frac{4}{1} - \frac{4}{3} + \frac{4}{5} - \frac{4}{7} + ...\)

OpenStudy (jiteshmeghwal9):

\[\Huge{\pi = {c \over d}=3.14159}\] 100+4 multiply by 8 add to 62000 this is approximately measure of the circumfrence of the circle whose diameter is 20,000 which mathematically can be written as :- \(\Huge{62832\over20000}=3.14\)

OpenStudy (jiteshmeghwal9):

@Callisto @UnkleRhaukus @ganeshie8 @lgbasallote @ParthKohli have a look :)

OpenStudy (anonymous):

is that ramanujan approximation

OpenStudy (jiteshmeghwal9):

no its Arraybhatta's @sara12345 :)

OpenStudy (jiteshmeghwal9):

@hartnn @hba @Hero @AccessDenied have a look :)

OpenStudy (jiteshmeghwal9):

@Preetha mam have a look :)

OpenStudy (jiteshmeghwal9):

@mayankdevnani have a look :)